# Irreversibility in scalar active turbulence: The role of topological defects

### v3

@@ 1 Introduction
In this paper, we examine the interplay between IEPR and topological defects in a continuum model of active turbulence. Inspired by previous works [51, 52, 53, 54, 55], our model describes the density dynamics of swimmers immersed in a momentum-conserving fluid. Recently, In analogy with the turbulence of active nematics [12], we show that vortical flows can spontaneously form despite the absence of inertia. By analytically deriving the IEPR, we demonstrate that the dominant contribution to irreversibility stems from vortices. We also show that the IEPR is a lower bound of the total dissipation derived from linear irreversible thermodynamics (LIT). Finally, we propose a mapping between the shear stress, which stirs the fluid, and an effective nematic field. Such a mapping leads to delineating singularities in the density profile which shape the surrounding vortices. In all, our results reveal how the symmetries of vertical flows, set by topological defects, determine the spatial distribution of IEPR.
The paper is structured as follows. In Sec. 2, we start by introducing our continuum theory, and reviewing the associated phenomenology: laminar, vortex, and turbulent states. We then show that the zeros of density gradients correspond to nematic-like singularities, and numerically evaluate the corresponding defect statistics. In Sec. 3, we explicitly relate the IEPR to the vorticity, and provide a numerical quantification in the noiseless regime. We also find a generic relatation between the IEPR and the total dissipation. We find that in steady state the IEPR is a lower bound to the total EPR. Finally, by computing the vorticity around defects, we examine in detail how the spatial distribution of IEPR is determined by the symmetries of defects.
@@ 2.1 Coupled dynamics of density and stream function
$\langle{\Gamma_{\alpha\beta}({\bf\it r},t)\Gamma_{\mu\nu}({\bf\it r}^{\prime},t^{\prime})}\rangle=(\delta_{\alpha\mu}\delta_{\beta\nu}+\delta_{\alpha\nu}\delta_{\beta\mu})\delta({\bf\it r}-{\bf\it r}^{\prime})(t-t^{\prime}).$ (5)
@@
$\eta\nabla^{2}v_{\alpha}=\partial_{\alpha}p-\partial_{\beta}\left(\Sigma_{\alpha\beta}^{P}+\Sigma_{\alpha\beta}^{A}+\sqrt{2\eta T}\Gamma_{\alpha\beta}\right),\quad\nabla\cdot{\bf\it v}=0.$ (7)
To eliminate the pressure contribution in Eq. (7), we define the stream function $\psi({\bf\it r},t)$ as
@@
$\langle{\Lambda_{\psi}({\bf\it r},t)\Lambda_{\psi}({\bf\it r}^{\prime},t^{\prime})}\rangle=\delta({\bf\it r}-{\bf\it r}^{\prime})(t-t^{\prime}).$ (10)
@@ 2.2 Laminar, vortex and turbulent states
We numerically simulate AMH [Eqs. 1 and 9] with the integration scheme detailed in B. The parameters are fixed to a=b=0.1-a=b=0.1, $\kappa=0.1$, and $\eta=1.67$, and we vary the activity parameter $\zeta<0$, namely, considering contractile activity. The extensile case ($\zeta>0$) does not have turbulent phases, so it is not relevant for our study which focuses on the turbulent, defect-rich phase.
@@
Figure 1: For Active Model H [Eqs. (1) and (9)], the swimmer density $\phi$ (top) and the scaled stream function $\tilde{\psi}=\frac{\eta}{\zeta+\kappa}\psi$ (bottom) feature three possible steady states for different levels of activity $\zeta$: (a) laminar flow, (b) vortex, and (c) spatio-temporally chaos (referred to as the turbulent state). The black lines with arrows show the velocity field lines. Parameters a=b=0.1-a=b=0.1, $\kappa=0.1$, and $\eta=1.67$. (d) Velocity spectrum $S(q)$ [Eq. (11)] as a function of wavenumber $q$ (scaled by the largest system size) for $\zeta=-1.1$.
@@ 2.3 Topological defects
Figure 2: (a) Density field $\phi$ in the turbulent state. The markers indicate the locations of $\frac{1}{2}$ (green circles) and $\frac{1}{2}$ (red triangles) topological defects. Panels (b,c): zoomed–in view of the four squares marked in panel (a): comparison between the location of the defects, the nematic director field (black lines) along with (b) the density gradient $|\nabla\phi|$ and (c) the local curvature $H$. (d) Density of defects $\rho_{\rm D}=\langle N_{\rm D}\rangle/L^{2}$, where $\langle N_{D}\rangle$ is the defect number averaged over time and realizations, and (e) defect lifetime $\tau_{\rm D}$, defined as the average time between pair creation and annihilation, as functions of the activity parameter $\zeta$. Same parameters as in Fig. 1.
@@
In short, the analogy between shear stress and nematic tensor means that singularities in the stress profile must have $\pm\frac{1}{2}$ charges. Moreover, such defects can actually be detected simply by analyzing the profile of swimmer density $\phi$, without any information on the surrounding flow.
@@ 3 Quantifying irreversibility from vorticity and topological defects
3.1 Breakdown of time-reversal symmetry
@@ 3.3 Local measurement of irreversibility: the role of topological defects
In the turbulent state, numerical measurements of the local IEPR $\dot{\sigma}$ [Eq. (16)] show that it is typically larger at interfaces than in the bulks, and also higher in regions of large vorticity [Fig. 4(a)]. Remarkably, the inhomogeneities in $\dot{\sigma}$ can be rationalized in terms of patterns and symmetries of topological defects of the nematic field $Q$ [Eq. (12)]. In the next paragraphs, we show that our analytical predictions reproduce the IEPR profiles observed in the simulations.
For isolated defects, the symmetries of the local nematic orientation (two-fold and six-fold symmetries for $\frac{1}{2}$ and $\frac{1}{2}$ defects, respectively) clearly affect the profile of $\dot{\sigma}$ [Figs. 4(b),(d)]. For a pair of $\frac{1}{2}$ defects, $\dot{\sigma}$ is locally maximum in between the defects [Fig. 4(f)]; for a pair of $\frac{1}{2}$ and $\frac{1}{2}$ defects, $\dot{\sigma}$ is locally higher close to the $\frac{1}{2}$ defect [Fig. 4(h)]. These observations suggest a route to predict the profile $\dot{\sigma}$ from the symmetries and relative orientations of defect pairs.
Figure 4: (a) Spatial distribution of the local IEPR from numerical simulation. The figure refers to the same turbulent state as in Figs. 1(a) and 2(a). The white dashed line represent the phase boundary $\phi=0$. Panels (b,d,f,h): zoomed–in views of isolated defect with charge (b) $\frac{1}{2}$ and (d) $\frac{1}{2}$ and defect pairs (f,h) marked resp. with orange, pink, green and red circles in panel (a). Panels (b-i): comparison between simulations and analytical predictions. We numerically evaluate the relative angle $\Phi$ and the distance $d$ between defects [Sec. 3.3.2] for a proper comparison of the local IEPR with analytics. Same parameters as in Fig. 1.
@@
The flow field around isolated defects can be found analytically by assuming that the density field $\phi$, which determines the local stress [Eq. (6)], relaxes faster than the stream function $\psi$, which determines the local flow. The stream function $\psi$ associated with the flow field $\bf\it v$} of an isolated defect was derived in Ref. [74]. Using polar coordinates ${\bf\it r}=(r,\varphi)$ centered at the defect core, the solution can be written as $\psi=\bar{\psi}+\psi_{0}$, where
@@
The IEPRs computed in [Eq. (31)] reproduce the angular symmetries observed in numerical simulations [Figs. 4(b)-(e)]. The orientation of the anisotropic $\frac{1}{2}$ defects is described by the polarization vector defined as $p_{\alpha}=-\partial_{\beta}Q_{\alpha\beta}({\bf\it r}={\bf\it r}_{+})$, where $\mathbb{Q}$ is the nematic field [Eq. (12)] and $\bf\it r$$_{+}$ is the position of the defect core [30]. We observe that $\dot{\sigma}$ has an approximate mirror symmetry across the direction of polarization where the vorticity is higher. For $\frac{1}{2}$ defects, the region of high IEPR has a six-fold symmetry, in line with the symmetry of vorticity. The six-lobed profile predicted analytically [Eq. (31b) and Fig. 4(e)] is not easily seen in the numerical simulation [Fig. 4(d)]. While Eq. (31b) predicts a maximum at the center of the defect, the numerical IEPR appears to be minimal. We ascribe this discrepancy to the fact that (i) the $\frac{1}{2}$ defects in the numerical simulation may not be isolated due to the surrounding $\frac{1}{2}$ defects, and (ii) the contribution of $\frac{1}{2}$ defects [Eq. (31a)] is much larger than that of $\frac{1}{2}$.
Our prediction of $\dot{\sigma}$ [Eq. (31)] assume no-slip boundary conditions, and is qualitatively independent from the choice of boundary conditions close to the defect core (i.e. far from the boundary layer). A different choice would only change our prediction quantitatively while preserving the angular symmetries [74]. In F, we demonstrate the existence of an inner region ($r<r_{C}$) where the IEPR is insensitive to the boundary conditions for all isolated $\pm\frac{1}{2}$ defects.
In short, the local IEPR around isolated defects is spatially inhomogeneous, with symmetries stemming from the vorticity of the backflow around defects: for the $\frac{1}{2}$ defect, it is mirror-symmetric across the direction of the polarization; for the $\frac{1}{2}$ defect, it features a six-fold rotational symmetry. The proximity of defects of either charges induces perturbations of these profiles. Generally, perturbations of the IEPR profile are more pronounced for $\frac{1}{2}$ defects.
@@
For a pair of defects with charges $(q_{1},q_{2})$, the configuration is described by two parameters: the separation $d$ between the defect cores, and the relative angle $\Phi$ between the defect orientations. The orientation of the nematic director field $\mathbb{Q}$ [Eq. (12)] reads
@@
Outside the defect cores, $\mathbb{Q}$ has a constant amplitude $S=\sqrt{Q_{11}^{2}+Q_{12}^{2}}$, so that the variations of $\mathbb{Q}$ are entirely determined by the orientation $\theta_{N}$. The velocity field $\bf\it v$} associated with the backflow of the defect pair follows by solving the Stokes equation [Eq. (7)]:
@@
Finally, the local IEPR follows from substituting the solution for the flow [Eq. (33)] into Eq. (16). Below, we examine three cases at a fixed distance ($d=R/2$) when changing the orientation $\Phi$. For all profiles, the local vorticity changes as $\omega\to-\omega$ under the transformation $\Phi\to\Phi+\pi$; given that $\dot{\sigma}$ is quadratic in $\omega$ [Eq. (16)], it is invariant under such a transformation, so we consider only the range $0\leq\Phi\leq\pi/2$.
@@
Let us start with a pair of $\frac{1}{2}$ defects [Fig. 5(a)]. At $\Phi=0$, the defects have collinear and opposite polarizations that are parallel to the segment that connects the cores; the vortices around each defect interfere negatively, yielding low $\dot{\sigma}$. At $\Phi=\pi/2$, the polarizations are opposite and perpendicular to the connecting segment; the vortices around each defect coalesce into a large vortex, yielding a high $\dot{\sigma}$. In contrast, a $\frac{1}{2}$ pair does not yield large $\dot{\sigma}$ as we rotate defects [Fig. 5(b)], since vortices do not coalesce as for a $\frac{1}{2}$ pair. For a $(+\frac{1}{2},-\frac{1}{2})$ pair [Fig. 5(c)], $\dot{\sigma}$ is highest at $\Phi=\pi/2$ as the left vortex of the $\frac{1}{2}$ does not interfere negatively with the vortices of the $\frac{1}{2}$ defect.
We compare our analytical predictions with some defect pairs spontaneously formed in the turbulent state. For a $\frac{1}{2}$ pair where $(d,\Phi)\simeq(0.45R,0.45\pi)$ [Fig. 4(f)], $\dot{\sigma}$ is higher between defects, in agreement with our predictions [Fig. 5(a)]. For a $(+\frac{1}{2},-\frac{1}{2})$ pair where $(d,\Phi)\simeq(0.49R,0.95\pi)$ [Fig. 4(h)], $\dot{\sigma}$ is higher close to the $\frac{1}{2}$ defect, and symmetric with respect to its polarization orientation, in agreement with our predictions [Fig. 5(c)]. In both cases, we attribute the slight quantitative deviations to the spatial asymmetries in the numerical solution of the nematic field.

### v4

# arXiv:2507.06073v4[cond-mat.stat-mech] 02 Mar 2026

@@ 1 Introduction
In this paper, we examine the interplay between IEPR and topological defects (TDs) in a continuum model of active turbulence. Inspired by previous works [51, 52, 53, 54, 55], our model describes the density dynamics of swimmers immersed in a momentum-conserving fluid. Recently, In analogy with the turbulence of active nematics [12], we show that vortical flows can spontaneously form despite the absence of inertia. By analytically deriving the IEPR, we demonstrate that the dominant contribution to irreversibility stems from vortices. We also show that the IEPR is a lower bound of the total dissipation derived from linear irreversible thermodynamics (LIT). Finally, we propose a mapping between the shear stress, which stirs the fluid, and an effective nematic field. Such a mapping leads to delineating singularities in the density profile which shape the surrounding vortices. In all, our results reveal how the symmetries of vertical flows, set by topological defects, determine the spatial distribution of IEPR.
The paper is structured as follows. In Sec. 2, we start by introducing our continuum theory, and reviewing the associated phenomenology: laminar, vortex, and turbulent states. We then show that the zeros of density gradients correspond to nematic-like singularities, and numerically evaluate the corresponding defect statistics. In Sec. 3, we explicitly relate the IEPR to the vorticity, and provide a numerical quantification in the noiseless regime. We also find a generic relation between the IEPR and the total dissipation. We find that in steady state the IEPR is a lower bound to the total EPR. Finally, by computing the vorticity around defects, we examine in detail how the spatial distribution of IEPR is determined by the symmetries of defects.
@@ 2.1 Coupled dynamics of density and stream function
$\langle{\Gamma_{\alpha\beta}({\bf\it r},t)\Gamma_{\mu\nu}({\bf\it r}^{\prime},t^{\prime})}\rangle=(\delta_{\alpha\mu}\delta_{\beta\nu}+\delta_{\alpha\nu}\delta_{\beta\mu})\delta({\bf\it r}-{\bf\it r}^{\prime})\delta(t-t^{\prime}).$ (5)
@@
$\eta\nabla^{2}v_{\alpha}=\partial_{\alpha}p-\partial_{\beta}\left(\Sigma_{\alpha\beta}^{P}+\Sigma_{\alpha\beta}^{A}+\sqrt{2\eta T}\Gamma_{\alpha\beta}\right),\quad\nabla\cdot{\bf\it v}=0\,$, (7)
where
$f^{A}_{\alpha}=\partial_{\beta}\Sigma_{\alpha\beta}^{A}$ (8)
is the active force and $f_{\alpha}=\partial_{\beta}(\Sigma^{P}_{\alpha\beta}+\Sigma^{A}_{\alpha\beta})$ is the total force. To eliminate the pressure contribution in Eq. (7), we define the stream function $\psi({\bf\it r},t)$ as
@@
$\langle{\Lambda_{\psi}({\bf\it r},t)\Lambda_{\psi}({\bf\it r}^{\prime},t^{\prime})}\rangle=\delta({\bf\it r}-{\bf\it r}^{\prime}){\delta}(t-t^{\prime}).$ (11)
@@ 2.2 Laminar, vortex and turbulent states
We numerically simulate AMH [Eqs. 1 and 10] with the integration scheme detailed in B. The parameters are fixed to a=b=0.1-a=b=0.1, $\kappa=0.1$, and $\eta=1.67$, and we vary the activity parameter $\zeta<0$, namely, considering contractile activity. The extensile case ($\zeta>0$) does not have turbulent phases, so it is not relevant for our study which focuses on the turbulent, defect-rich phase.
@@
Figure 1: For Active Model H [Eqs. (1) and (10)], the swimmer density $\phi$ (top) and the scaled stream function $\tilde{\psi}=\frac{\eta}{\zeta+\kappa}\psi$ (bottom) feature three possible steady states for different levels of activity $\zeta$: (a) laminar flow, (b) vortex, and (c) spatio-temporal chaos (referred to as the turbulent state). The black lines with arrows show the velocity field lines. Parameters a=b=0.1-a=b=0.1, $\kappa=0.1$, and $\eta=1.67$. (d) Velocity spectrum $S(q)$ [Eq. (16)] as a function of wavenumber $q$ (scaled by the largest system size) for $\zeta=-1.1$.
@@ 2.3 Topological defects
$S=\sqrt{Q_{11}^{2}+Q_{12}^{2}}=\frac{1}{2}|\nabla\phi|^{2}$ (16)
which vanishes at the critical points of $\phi$, namely at the center of the TDs, as expected. In active nematics, the core radius is the typical length scale in which $S$ varies from 1 to 0; however, here $S$ is not an order parameter field. The core is set by the typical spatial variation of $|\nabla\phi|$ around the critical points of $\phi$, see for instance Fig. 2(b). We refer the reader to F for further discussion on the core size.
Figure 2: (a) Density field $\phi$ in the turbulent state ($\zeta=-0.16$). The markers indicate the locations of $\frac{1}{2}$ (green circles) and $\frac{1}{2}$ (red triangles) topological defects. Panels (b,c): zoomed–in view of the four squares marked in panel (a): comparison between the location of the defects, the nematic director field (black lines) along with (b) the density gradient $|\nabla\phi|$ and (c) the local curvature $H$. (d) Density of defects $\rho_{\rm D}=\langle N_{\rm D}\rangle/L^{2}$, where $\langle N_{D}\rangle$ is the defect number averaged over time and realizations, and (e) defect lifetime $\tau_{\rm D}$, defined as the average time between pair creation and annihilation, as functions of the activity parameter $\zeta$. Parameters a=b=0.1-a=b=0.1, $\kappa=0.1$, and $\eta=1.67$.
@@
In short, the analogy between shear stress and nematic tensor means that singularities in the stress profile must have $\pm\frac{1}{2}$ charges. Moreover, such defects can actually be detected simply by analyzing the profile of swimmer density $\phi$, without any information on the surrounding flow. In what follows, to analyze the turbulent phase of AMH, we use three equivalent levels of description: (i) the density of microswimmers measured by the scalar field $\phi$ [Fig. 2(a)], (ii) the stresses (passive and active) represented by the nematic order parameter $\mathbb{Q}$ [Eq. (13) and Fig. 3(a)], and (iii) the forces (passive and active, see Eq. (8)) which are proportional to $\nabla\cdot\mathbb{Q}$ [Fig. 3(b)]. The forces are the sources of the active flows [Eq. (7)] which stir the topological defects.
Figure 3: (a) Nematic order parameter $\mathbb{Q}$ [Eqs. (13) and (14)] extracted from the density field $\phi$ in Fig. 2(a). The black lines correspond to the director field representing the nematic orientation $\theta_{N}$ [Eq. (15)]. The blue colorscale corresponds to the amplitude $S$ of $\mathbb{Q}$ [Eq. (16)] that is maximal at interfaces between dense and dilute regions [Fig. 2(a)]. (b) Total force field is proportional to $\nabla\cdot\mathbb{Q}$ [Eq. (8)]. The gray colorscale corresponds to the amplitude of the total force that is localized at the interfaces. The colored arrows refer to the orientation of the force vectors, see inset.
@@ 3 Quantifying irreversibility from vorticity and topological defects
3.1 IEPR as a measure of irreversibility
@@ 3.3 Local measurement of irreversibility: the role of topological defects
In the turbulent state, the local IEPR $\dot{\sigma}$ fluctuates in the form of sharply localized regions that continuously appear and disappear in time; we refer the reader to the Supplementary video. We report in Fig. 5(a) a typical instantaneous configuration, where the field $\dot{\sigma}$ is directly evaluated from $\psi$ [Eq. (19)], showing that $\dot{\sigma}$ is typically larger at interfaces (between small- and large-$\phi$ regions) than in the bulks, and also higher in regions of large vorticity. We observe two necessary conditions for regions of high IEPR to emerge: (i) the interfaces form S-shaped meandering curves, and (ii) the curved interfaces enclose pairs of $\frac{1}{2}$ TDs which are oriented perpendicular to the line connecting them [red circle in Fig. 5(a) and Figs. 5(d,e)]. In general, other configurations of paired TDs can produce a high IEPR locally: in a pair of $\frac{1}{2}$ and $\frac{1}{2}$ TDs, $\dot{\sigma}$ is locally higher close to the $\frac{1}{2}$ defect [Fig. 5(f)].
In what follows, we provide analytical predictions that reproduce the IEPR profiles observed in the simulations, and argue that such profiles can be rationalized in terms of the symmetries of the TDs in the nematic field $Q$ [Eq. (13)]. To this end, we examine idealized TDs by considering either isolated TDs of charge $\frac{1}{2}$ and $\frac{1}{2}$ [Sec. 3.3.1], or pairs of TDs obtained by superposing the orientation fields of two isolated defects [Sec. 3.3.2].
Figure 5: (a) Spatial distribution of the local IEPR $\dot{\sigma}$ [Eq. (19)] extracted from the same turbulent state as in Figs. 1(c) and 2(a). The white dashed lines represent the phase boundary $\phi=0$. The green dots and red triangles respectively refer to $\frac{1}{2}$ and $\frac{1}{2}$ topological defects. (b,c) Analytic predictions of $\dot{\sigma}$ around isolated topological defects [Eq. (32)] of charge $\frac{1}{2}$ [panel (b)] and $\frac{1}{2}$ [panel (c)]. White lines refer to the nematic orientation. (d-g) Comparison between simulations and analytical predictions for defect pairs. A pair with charges $(+\frac{1}{2},+\frac{1}{2})$ (d,e) and a pair with charges $(+\frac{1}{2},-\frac{1}{2})$ (f,g) marked respectively with red and green circles in panel (a). We numerically evaluate the relative angle $\Phi$ and the distance $d$ between defects [Sec. 3.3.2]. White lines refer to the nematic orientation.
@@
We predict analytically the intensity and the spatial symmetries of the local IEPR $\dot{\sigma}$ around isolated defects using Eq. (19). The flow field around isolated defects can be found analytically by assuming that (i) the amplitude of the nematic $Q$-tensor is constant, and (ii) the density field $\phi$, which determines the local stress [Eq. (6)], relaxes faster than the stream function $\psi$, which determines the local flow. The stream function $\psi$ associated with the flow field $\bf\it v$} of an isolated defect was derived in Ref. [74]. Using polar coordinates ${\bf\it r}=(r,\varphi)$ centered at the defect core, the solution can be written as $\psi=\bar{\psi}+\psi_{0}$, where
@@
The spatial symmetries of the stream function are reflected in the profile of the IEPR [Figs. 5(b,c)]. Our prediction of $\dot{\sigma}$ [Eq. (33)] assume no-slip boundary conditions, and is qualitatively independent of the choice of boundary conditions close to the defect core (i.e. far from the boundary layer). A different choice would only change our prediction quantitatively while preserving the angular symmetries [74]. In F, we demonstrate the existence of an inner region ($r<r_{C}$) where the IEPR is insensitive to the boundary conditions for all isolated $\pm\frac{1}{2}$ defects.
In short, the local IEPR around isolated defects is spatially inhomogeneous, with symmetries stemming from the vorticity of the backflow around defects: for the $\frac{1}{2}$ defect, it is mirror-symmetric across the direction of the polarization; for the $\frac{1}{2}$ defect, it features a six-fold rotational symmetry. The proximity of defects of opposite charges induces perturbations of these profiles. Generally, perturbations of the IEPR profile are more pronounced for $\frac{1}{2}$ defects. In typical configurations [Fig. 5(a)], there are no rigorously isolated defects due to the highly correlated distortions of the interfaces across the sample.
The assumption of a constant amplitude of the $Q$-tensor that underlies our derivation leads to quantitative differences between the analytic and the simulated profiles. In simulations, the active force $f^{A}_{\alpha}\propto\partial_{\beta}Q_{\alpha\beta}$ [Eq. (8)] around $\pm\frac{1}{2}$ defects is localized along the interfaces of $\phi$, and vanishes at the center of the defects, whereas it is maximal at the center in our analytic solution. The analysis and justification of these discrepancies are deferred to G.
@@
To predict the profiles of the local IEPR $\dot{\sigma}$ around defect pairs, we construct the director field $\theta_{N}({\bf\it r})$ by superposing the textures of isolated defects. For a pair of defects with charges $(q_{1},q_{2})$, the configuration is described by two parameters: the separation $d$ between the defect cores, and the relative angle $\Phi$ between the defect orientations. The orientation of the nematic director field [Eq. (13)] reads
@@
We assume that the scale of the defect cores is small compared to the typical variations of the IEPR profile, and that the amplitude $S=\sqrt{Q_{11}^{2}+Q_{12}^{2}}$ is constant outside the defect cores, so that the variations of $\mathbb{Q}$ are entirely determined by the orientation $\theta_{N}$. The velocity field $\bf\it v$} associated with the backflow of the defect pair follows by solving the Stokes equation [Eq. (7)] by using a Green function: $v_{\alpha}({\bf\it r})=\int\mathrm{d}^{2}{\bf\it r}{{}^{\prime}}G_{\alpha\beta}({\bf\it r}-{\bf\it r}^{\prime})f_{\beta}({\bf\it r}^{\prime})$ where $f$ is the force acting on the fluid. Using $f_{\beta}=\partial_{\gamma}\Sigma_{\beta\gamma}$ we can write:
@@
Finally, the local IEPR follows from substituting the solution for the flow [Eq. (35)] into Eq. (18). Below, we examine three cases at a fixed distance ($d=R/2$) when changing the orientation $\Phi$. The orientation of the anisotropic $\frac{1}{2}$ defects is described by the polarization vector defined as $p_{\alpha}=-\partial_{\beta}Q_{\alpha\beta}({\bf\it r}={\bf\it r}_{+})$, where $\mathbb{Q}$ is the nematic field [Eq. (13)] and $\bf\it r$$_{+}$ is the position of the defect core [30]. For all profiles, the local vorticity changes as $\omega\to-\omega$ under the transformation $\Phi\to\Phi+\pi$; given that $\dot{\sigma}$ is quadratic in $\omega$ [Eq. (18)], it is invariant under such a transformation, so we consider only the range $0\leq\Phi\leq\pi/2$.
@@
Let us start with a pair of $\frac{1}{2}$ defects [Fig. 6(a)]. At $\Phi=0$, the defects have collinear and opposite polarizations that are parallel to the segment that connects the cores; the vortices around each defect interfere negatively, yielding low $\dot{\sigma}$. At $\Phi=\pi/2$, the polarizations are opposite and perpendicular to the connecting segment; the vortices around each defect coalesce into a large vortex, yielding a high $\dot{\sigma}$. In contrast, a $\frac{1}{2}$ pair does not yield large $\dot{\sigma}$ as we rotate defects [Fig. 6(b)], since vortices do not coalesce as for a $\frac{1}{2}$ pair. For a $(+\frac{1}{2},-\frac{1}{2})$ pair [Fig. 6(c)], $\dot{\sigma}$ is highest at $\Phi=\pi/2$ as the left vortex of the $\frac{1}{2}$ does not interfere negatively with the vortices of the $\frac{1}{2}$ defect. In short, the vortices generated by the total force regulate the localization of IEPR. Since the total force is localized along the interfaces between regions of small and large $\phi$ [Fig. 3(b)], it follows that $\dot{\sigma}$ is also highest at interfaces.
We compare our analytical predictions with some defect pairs spontaneously formed in the turbulent state. For a $\frac{1}{2}$ pair where $(d,\Phi)\simeq(0.45R,0.45\pi)$ [Fig. 5(e)], $\dot{\sigma}$ is higher between defects, in agreement with our predictions [Fig. 6(a)]. For a $(+\frac{1}{2},-\frac{1}{2})$ pair where $(d,\Phi)\simeq(0.49R,0.95\pi)$ [Fig. 5(g)], $\dot{\sigma}$ is higher close to the $\frac{1}{2}$ defect, and symmetric with respect to its polarization orientation, in agreement with our predictions [Fig. 6(c)]. In both cases, we attribute the slight quantitative deviations to the spatial asymmetries in the numerical solution of the nematic field.